OpenStudy (anonymous):

I GIVE MEDALS Use the following equation: f(x) = x3 – 2x2 – 2x + 1 According to the Rational Root Theorem, which of the following would be a factor of the function? Select one: a. x – 1 b. x + 1 c. x2 – 1 d. x2 + 1

OpenStudy (gamer56):

What do you think it is?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

what do you game on?

OpenStudy (gamer56):

!That is what id put @chantz417 ! iF i helped you then please fan and give me a medal!

OpenStudy (dumbcow):

only possible rational roots are 1 or -1 plug in each to see which one works

OpenStudy (dumbcow):

if x=1 works then factor is " x-1" if x =-1 works then factor is "x+1" if both work then factor is "x^2 -1" --> x^2-1 = (x-1)(x+1)

OpenStudy (anonymous):

plug it in for x right?

OpenStudy (dumbcow):

yes and see if you get y=0

OpenStudy (anonymous):

-1 works

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

yes correct :)

OpenStudy (anonymous):

thank you